Differentiate the Logarithmic functions:

1. 𝑦 = π‘₯^3𝑙𝑛^2π‘₯ + π‘₯^2𝑙𝑛π‘₯^3 + 1

2. 𝑔(π‘₯) = 𝑙𝑛 (π‘₯^2+1/π‘₯^3+5)

3. β„Ž(π‘₯) = π‘™π‘œπ‘”(π‘π‘œπ‘‘4π‘₯ βˆ’ 𝑐𝑠𝑐4π‘₯)

4. 𝑦 = 𝑙𝑛(π‘₯^3𝑠𝑖𝑛2π‘₯)

5. 𝑦 = 𝑙𝑛(π‘π‘œπ‘ 5π‘₯ + 𝑠𝑖𝑛5π‘₯)



1
Expert's answer
2020-12-20T18:01:55-0500

1.

"y'=(𝑥^3 \ln^2x + 𝑥^2\ln x^3+ 1)'"

"=3x^2 \ln^2x+x^3(2\ln x)(\dfrac{1}{x})+2x(3\ln x)+3x^2(\dfrac{1}{x})+0"

"=3x^2 \ln^2x+2x^2\ln x+6x\ln x+6, x>0"

2.


"𝑔'(𝑥) = (\ln (𝑥^2+1/𝑥^3+5))'"

"=\dfrac{1}{x^2+\dfrac{1}{x^3}+5}(2x-\dfrac{3}{x^4 }+0)"

"=\dfrac{2x^5-3}{x^6+5x^4+x}"

3.


"ℎ'(𝑥) = (\log(\cot4𝑥 − \csc4𝑥))'"

"= \dfrac{1}{\ln 10}\cdot\dfrac{4(-\dfrac{1}{\sin^24x})-4(-\dfrac{\cos4x}{\sin^24x})}{\cot4𝑥 − \csc4𝑥}"

"= \dfrac{4}{\ln 10}\cdot\dfrac{-1+\cos 4x}{\sin 4x(\cos4𝑥 −1)}"

"= \dfrac{4}{\ln 10}\csc 4x"

"\cot4𝑥 − \csc4𝑥>0=>\dfrac{\cos 4x-1}{\sin 4x}>0"

"=>\dfrac{1-\cos 4x}{\sin 4x}<0=>\sin 4x<0"


4.


"y' = (\ln(𝑥^3\sin 2𝑥))'="

"=\dfrac{3x^2\sin 2x+2x^3\cos 2x}{x^3\sin 2x}"

"=\dfrac{3}{x}+2\cot 2x, \ x\sin 2x>0"

5.


"y' =(\ln(\cos5𝑥 + \sin5𝑥))'"

"=\dfrac{1}{\cos5𝑥 + \sin5𝑥}(-5\sin 5x+5\cos 5x)"

"=5\cdot\dfrac{\cos 5x-\sin 5x}{\cos5𝑥 + \sin5𝑥},\ \cos5𝑥 + \sin5𝑥>0"